# SeMaScore : a new evaluation metric for automatic speech recognition tasks

### v1

@@ Abstract
Index Terms—  Evaluation metric, ASR, Semantic Similarity, Error rate
@@ 1 Introduction and motivation
Over the years, numerous research works have sought to provide alternatives to WER [4, 5] as an evaluation metric and have demonstrated its limited alignment with human judgment when applied in the context of spoken conversations [6]. Recent studies [7, 8, 9], have introduced the utilization of pre-trained BERT-based [10] models to compute a semantic distance metric between the ground truth and hypothesis. However, semantic distance metric overly prioritizes some words over others making it unreliable. In another study presented in [11], the authors employed BERTScore [12] as an evaluation criterion for ASR systems. However, BERTscore has some limitations as well and this is presented in Table 1. Table 1 presents two cases where BERTscore fails to provide an accurate score. The first example illustrates two entirely mismatched words, while the second example showcases a sentence with repetitions—a common mistake in ASR due to encoder-decoder model hallucination issues. In both instances, BERTscore assigns a higher similarity score and fails to detect the errors, despite the sentences being flawed. Motivated by these insights, we created a new metric which focuses on both the error rate and a more robust semantic similarity score for the comprehensive evaluation of ASR models.
@@ 2.1 Segment mapping
ASR evaluation, being a sequential task, does not require the complex token-level mapping used in BERTscore. In our approach, we adopt a sentence alignment technique based on the Levenshtein distance to create a mapping between GT and H. To establish this mapping, we first compute the character-level edit distance between GT and H. We then attempt to group the sentences at the word level when possible; otherwise, we group words into segments for sentence-level mapping. Table 2 provides an example of this mapping process. As demonstrated in the table, our approach accommodates cases where words are split (e.g., ‘sandwich’ $\rightarrow$ ‘sand wich’), combined (e.g., ‘have a’ $\rightarrow$ ‘havea’), or misrecognized (e.g., ‘have’ $\rightarrow$ ‘hav’). This adaptable mapping strategy allows us to precisely identify the segment-wise correspondence ($GT_{M}$ and $H_{M}$), between GT and H.
@@ 2.2 Computing segment score
As explained in Section 1, by limiting to cosine similarity between token-level embeddings, as done in BERTscore, has its disadvantages of assigning a higher score. Therefore, we propose the introduction of a penalty for the cosine similarity score by multiplying it with (1 - match error rate (MER)). We leverage MER, an error-based metric used to assess character-level mistakes. MER is used instead of the traditional CER, as MER is bounded between 0 and 1.
@@
(a) BERTscore
(b) SeMaScore
Fig. 1: Distribution of various metrics for disordered speech
(a) BERTscore
(b) SeMaScore
Fig. 2: Distribution of various metrics for noisy speech
@@ 2.4 Metric calculation
3 ASR Analysis and results
We discuss the diverse settings employed to evaluate our SeMaScore metric. We outline the multiple datasets utilized in our analysis and describe the comparison methodology and results. All results were evaluated against the current SOTA BERTscore. All contextual embeddings were obtained using pre-trained deberta-large-mnli model. Revisiting the examples in Table 1 where BERTscore fails, we can see that SeMaScore evaluated the hypotheses better by penalizing them appropriately. All the codes, datasets employed, and extended results are available from the following link: link
3.1 Disordered speech
Fig.1 presents the results obtained when BERTscore and SeMaScore are used to evaluate the disordered speech. It is clear from the figures that both BERTscore and SeMaScore correlate with human assessments. From figures 0(a) and 0(b), we can infer that for category 2, BERTscore has a mean of 0.9 with a standard deviation of 0.05 and SeMaScore has a mean of 0.68 with a standard deviation of 0.16 respectively. Since category 2 contains hypotheses whose meaning is different from that of the ground truths, we want the metrics to give a lower score which is well satisfied by SeMaScore.
@@ 3.2 Noisy speech
Fig.2 portrays the results discovered when BERTscore and SeMaScore are used to evaluate the noisy speech. The figures show that SeMaScore correlate well with SNR levels. We observe that, at lower SNR level 2.5, the hypotheses generated correspond to assessment category 2. From figures 1(a) and 1(b), we can infer that for SNR level 2.5, BERTscore has a mean of 0.9 with a standard deviation of 0.06 and SeMaScore has a mean of 0.71 with a standard deviation of 0.22 respectively. Since SNR level 2.5 includes hypotheses whose meaning differs from that of the ground truth, we expect the metrics to produce a lower score, which is again effectively delivered by SeMaScore.

### v2

@@ Abstract
keywords:
Evaluation metric, Speech recognition, Semantic Similarity, Error rate
@@ 1 Introduction and motivation
Over the years, numerous research works have sought to provide alternatives to WER [4, 5] as an evaluation metric and have demonstrated its limited alignment with human judgment when applied in the context of spoken conversations [6]. Recent studies [7, 8, 9, 10] have introduced the utilization of pre-trained BERT-based [11] models to compute a semantic distance metric between the ground truth and hypothesis. However, the semantic distance metric calculated at sentence-level overly prioritizes some words over others, making it unreliable. In BERTScore [12], the authors calculated the semantic distance at token-level as an evaluation criterion for ASR systems. However, BERTScore has some limitations as well, and this is presented in Table 1. Table 1 presents two cases where BERTScore fails to provide an accurate score. The first example illustrates two entirely mismatched words, while the second example showcases a sentence with repetitions—a common mistake in ASR due to encoder-decoder model hallucination issues. In both instances, BERTScore assigns a higher similarity score and fails to detect the errors, despite the sentences being flawed. The authors identified this nature of BERTScore and proposed baseline rescaling, which tones down the value of BERTScore. However, we observed that some of the examples still fail after rescaling. For instance, if we consider the same example 1 of Table 1, even after rescaling, the BERTScore drops marginally from 0.98 to 0.86. Further, baseline rescaling involves the computation of BERTScore over a corpus of million sentences for every sub-task in every language. Motivated by these insights, we created a new metric that focuses on both the error rate and a more robust semantic similarity score for the comprehensive evaluation of ASR models.
@@ 2.1 Segment mapping
ASR evaluation, being a sequential task, does not require the complex token-level mapping used in BERTScore. In our approach, we adopt a sentence alignment technique based on the Levenshtein distance to create a mapping between GT and H. To find the mapping between the sentences, we first align GT and H based on the character-level edit distance. Subsequently, we identify corresponding words or groups of words (segments) between GT and H that align with each other. Table 2 provides an example of this segment mapping approach, where aligned portions of sentences correspond to single words or groups of words. As demonstrated in the table, our approach accommodates cases where words are split (e.g., ‘sandwich’ $\rightarrow$ ‘sand wich’), combined (e.g., ‘have a’ $\rightarrow$ ‘havea’), or misrecognized (e.g., ‘have’ $\rightarrow$ ‘hav’). This adaptable mapping strategy allows us to precisely identify the segment-wise correspondence ($GT_{M}$ and $H_{M}$), between GT and H.
@@ 2.2 Computing segment score
As explained in Section 1, by limiting to cosine similarity between token-level embeddings, as done in BERTScore, has its disadvantages of assigning a higher score. Therefore, we propose the introduction of a penalty for the cosine similarity score by multiplying it with (1 - match error rate (MER)). We leverage MER [6], an error-based metric used to assess character-level mistakes. MER is used instead of the traditional CER, as MER is bounded between 0 and 1.
@@ 2.4 Metric calculation
3 ASR Analysis and results
We discuss the diverse settings employed to evaluate our SeMaScore metric. We outline the multiple datasets utilized in our analysis and describe the comparison methodology and results. All results were evaluated against the current SOTA BERTScore. All contextual embeddings were obtained using pre-trained roberta-base model. Revisiting the examples in Table 1 where BERTScore fails, we can see that SeMaScore evaluated the hypotheses better by penalizing them appropriately. All the codes, datasets employed, and results are available from the following link: https://github.com/zenlab-edgeASR/SeMaScore.
(a) BERTScore
(b) SeMaScore
Figure 1: Distribution of various metrics for disordered speech
(a) Category 0
(b) Category 1
(c) Category 2
Figure 2: Correlation between SeMaScore and WER for disordered speech
(a) Category 0
(b) Category 1
(c) Category 2
Figure 3: Correlation between SeMaScore and BERTScore for disordered speech
3.1 Disordered speech
Figure 1 presents the results obtained when BERTScore and SeMaScore are used to evaluate the disordered speech. It is clear from the figures that both BERTScore and SeMaScore correlate with human assessments. From figures 1(a) and 1(b), we can infer that for category 2, BERTScore has a mean of 0.9 with a standard deviation of 0.05 and SeMaScore has a mean of 0.68 with a standard deviation of 0.16 respectively. Since category 2 contains hypotheses whose meaning is different from that of the ground truths, we want the metrics to give a lower score which is well satisfied by SeMaScore.
Figure 2 and Figure 3 depict scatter plots illustrating the correlation between SeMaScore and both WER and BERTScore. The plots show a negative correlation between SeMaScore and WER, as well as a positive correlation between SeMaScore and BERTScore across all categories. From Figure 2, it is apparent that the range of SeMaScore values corresponding to a specific WER value has a wide spread, indicating the consideration of semantic content. Figure 3(c) demonstrates a strong correlation between our metric and BERTScore. However, SeMaScore produces a sensible range of scores (0.3 to 1) for categories containing semantically diverse sentences, compared to BERTScore (0.85 to 1).
(a) BERTScore
(b) SeMaScore
Figure 4: Distribution of various metrics for noisy speech
(a) SNR level 2.5
(b) SNR level 7.5
(c) SNR level 12.5
Figure 5: Correlation between SeMaScore and WER for noisy speech
(a) SNR level 2.5
(b) SNR level 7.5
(c) SNR level 12.5
Figure 6: Correlation between SeMaScore and BERTScore for noisy speech
@@ 3.2 Noisy speech
Figure 4 portrays the results discovered when BERTScore and SeMaScore are used to evaluate the noisy speech. The figures show that SeMaScore correlates well with SNR levels. We observe that, at a lower SNR level 2.5, the hypotheses generated correspond to assessment category 2. From figures 4(a) and 4(b), we can infer that for SNR level 2.5, BERTScore has a mean of 0.9 with a standard deviation of 0.06 and SeMaScore has a mean of 0.71 with a standard deviation of 0.22 respectively. Since SNR level 2.5 includes hypotheses whose meaning differs from that of the ground truth, we expect the metrics to produce a lower score, which is again effectively delivered by SeMaScore.
Figure 5 and Figure 6 present scatter plots illustrating the correlation between SeMaScore with WER and BERTScore across various SNR levels. In Figure 5(a), at a lower SNR level of 2.5, where the hypotheses may contain more errors, SeMaScore values exhibit a broader distribution similar to WER. In Figure 5(b) and Figure 5(c) at higher SNR levels, both WER and SeMaScore values are confined to a lower range, as expected. Overall, WER and SeMaScore correlate well for this task. Similar to our findings in disordered speech experiments, Figure 6 showcases a strong correlation between our metric and BERTScore. SeMaScore produces scores within a broad range (0.1 to 1), in contrast to the higher BERTScore values (0.75 to 1).